OpenStudy (lizz123):

Medal: Examine the table of values. (under the second question) 1. Write an equation that gives the value of y for any given value of x. 2. State the slope and y-intercept of the line that your equation represents. Table: x −3 −2 −1 0 1 2 3 y 15 10 5 0 5 10 15

OpenStudy (lizz123):

@Nnesha

OpenStudy (anonymous):

y= Mx + 0 put in any value and M = -5. Gives us Y= -5X

OpenStudy (lizz123):

So would my answer be -5x=-5+0

OpenStudy (lizz123):

@ArsalanAijaz

OpenStudy (lizz123):

@dan815

OpenStudy (anonymous):

Standard form of linear equation is Y=mx+C. C is the y-intercept (where x = 0), that's clearly 0 from the data so our equation is reduced to Y=mx. Plug in any value of X and Y and you get M

OpenStudy (lizz123):

2. (0,0)

OpenStudy (anonymous):

Using 0,0 is not a good idea cuz we have already used it for Y-intercept and secondly M could be anything then cuz 0 multiplied by whatever is 0

OpenStudy (lizz123):

0,5

OpenStudy (lizz123):

@ArsalanAijaz

OpenStudy (anonymous):

wait

OpenStudy (lizz123):

k

OpenStudy (anonymous):

can you post a screenshot?

OpenStudy (lizz123):

i dont know how to do tht

OpenStudy (anonymous):

Y= -5X. It's the same question!

OpenStudy (lizz123):

@ArsalanAijaz